OpenStudy (anonymous):

simplify. 3-7x over x+4 + x+8 over x + 4 = ?

jhonyy9 (jhonyy9):

(3-7x)/(x+4) +(x+8)/(x+4)=(3-7x+x+8)/(x+4)=(11-6x)/(x+4)